# Protap: A Benchmark for Protein Modeling on Realistic Downstream Applications

### v2

@@ Abstract
Recently, extensive deep learning architectures and pretraining strategies have been explored to support downstream protein applications. Additionally, domain-specific models incorporating biological knowledge have been developed to enhance performance in specialized tasks. In this work, we introduce Protap, a comprehensive benchmark that systematically compares backbone architectures, pretraining strategies, and domain-specific models across diverse and realistic downstream protein applications. Specifically, Protap covers five applications: three general tasks and two novel specialized tasks, i.e., enzyme-catalyzed protein cleavage site prediction and targeted protein degradation, which are industrially relevant yet missing from existing benchmarks. For each application, Protap compares various domain-specific models and general architectures under multiple pretraining settings. Our empirical studies imply that: (i) Though large-scale pretraining encoders achieve great results, they often underperform supervised encoders trained on small downstream training sets. (ii) Incorporating structural information during downstream fine-tuning can match or even outperform protein language models pretrained on large-scale sequence corpora. (iii) Domain-specific biological priors can enhance performance on specialized downstream tasks. Code and datasets are publicly available at https://github.com/Trust-App-AI-Lab/protap.
1 Introduction
@@
Proteins serve as the central executors of biological activities, regulating a wide range of critical biological processes through their complex three-dimensional structures and dynamic properties. A precise understanding of protein function and interactions is critical across many applications [1, 2]. For instance, directed evolution of enzymes can endow proteins with novel functions [3]. Accurate prediction of protein–ligand interactions (PLIs) can largely accelerate drug discovery [4, 5, 6]. These application areas underscore the immense potential of deep learning in protein analysis.
@@
Various deep learning approaches leveraging protein sequences and structures have been developed for protein-related applications. For instance, general sequence models such as LSTM [7] and Transformer [8] have been employed to extract amino acid sequence patterns for tasks like stability landscape prediction [9]. Geometric Graph Neural Networks (GNNs), including GVP-GNN [10] and EGNN [11], have demonstrated remarkable effectiveness in modeling 3D molecular structures. Motivated by these successes, geometric GNNs have been further applied to structure-based protein modeling tasks [12, 13]. Recently, Graphformer [14] and Transformer-M [15] have incorporated structural biases into transformer-based architectures. These emerging sequence-structure hybrid models show promising ability in the general protein representation learning.
Recently, the remarkable success of large-scale pretraining models in image and text has inspired similar advances in protein modeling. Masked language modeling has been extended to predict the masked amino acids in protein sequences, resulting in protein language models like ESM and ProteinBERT [16, 17, 18]. GearNet [19] explores the multi-view contrastive learning on protein structures. Additionally, some methods, such as OntoProtein [20], leverage functional annotations as supervision signals for pretraining.
@@
Apart from these general model architectures and pretraining tasks for protein modeling, significant progress has been made in developing domain-specific models tailored for various realistic downstream protein applications. For instance, protein function prediction models such as DPFunc [21] leverage sequence, structural, and domain-level information to enhance the accuracy of predicting the Gene Ontology annotations of proteins. Additionally, to improve enzyme modeling, UniZyme [22] integrates the energy frustration matrix and enzyme active-site knowledge into a transformer-based framework. In the domain of proteolysis-targeting chimera (PROTAC) modeling, dedicated methods such as DeepProtacs [23] and ETProtacs [24] have been developed to accurately capture the ternary interactions essential for targeted protein degradation.
@@
Given the extensive variety of general model architectures, pretraining strategies, and domain-specific models, there remains a gap in systematically benchmarking general pretrained models alongside domain-specific models. With numerous architectures and pretraining strategies tailored for real-world downstream tasks, a natural question arises: do existing architectures and strategies exhibit distinct advantages across specific protein applications? However, current benchmarks predominantly focus on specific pretrained model categories, such as protein language models in [25, 26] and geometric GNNs in [27]. As a result, they lack a comprehensive evaluation of general protein model architectures, pretraining strategies, and domain-specific models across realistic biological applications. To address these gaps, we introduce Protap, a standardized benchmark that systematically compares backbone architectures, pre-training strategies, and domain models on diverse realistic downstream applications. Our main contributions are:
@@
  • We identify and integrate realistic applications from existing literature and databases to support comprehensive evaluations. In addition to three general applications, Protap introduces two novel specialized applications: enzyme-catalyzed protein cleavage site prediction and targeted protein degradation by PROTACs, which are biological processes not covered by prior benchmarks.
@@
  • The proteins covered in the benchmark evaluation are comprehensive and diverse, which cover enzymes, receptors, drugs, etc. The tasks are diverse, which include single protein modeling (function, mutation), interaction modeling such as PLI, enzyme-substrate modeling, and complex interaction process modeling (PROTACs).
@@
  • We compare a large number of protein pretraining models and domain models on five protein applications. This offers insights into the development of protein foundation models and the design of domain-specific models for downstream applications.
Table 1: Comparison of benchmark coverage for protein modeling. indicates the presence of this dimension, while ○ denotes its absence. We compare across three dimensions: applications (specialized vs. general), pretraining tasks, and model architectures (domain-specific vs. general).
Our Protap PEER [25] ProteinWorkshop [27] TAPE [9] ProteinGLUE [26]
Targeted protein degradation by Proteolysis-targeting chimeras
General Applications Protein–ligand interactions
Function prediction
Mutation effect prediction
Pretraining Tasks Masked language modeling
Multi-view contrastive learning
Protein family prediction
General Architectures Protein language models
Geometric GNNs
Sequence-structure hybrid models
Domain-specific Architectures Protein function models
Enzyme domain models
Ternary complexes models
2 Related Works
@@
Existing benchmarks for protein modeling primarily emphasize sequence-based models. TAPE [9], one of the earliest benchmarks, introduces standardized sequence-centric tasks including structure prediction and remote homology detection. PEER [25] expands the sequence-based scope to cover protein–protein interactions and functional annotations. ProteinGLUE [26] is another sequence-based benchmark designed to evaluate sequence pretraining methods. ProteinGym [28] establishes an evaluation framework for mutation effect prediction with protein language models. ProteinBench [29] evaluates protein foundation models mainly in protein generation tasks. Among existing benchmarks, ProteinWorkshop [27] incorporates the equivariant graph neural networks to evaluate the structure-based models. However, ProteinWorkshop lacks a unified benchmarking pipeline that systematically compares sequence-based, structure-based, and hybrid models within a standardized framework.
@@
Furthermore, existing benchmarks predominantly focus on general model architectures such as LSTMs [7], Transformers [8], and EGNNs [11]. In contrast, Protap explores the advantages of domain-specific model designs and the integration of domain knowledge. Our Protap systematically benchmarks domain-specific models alongside pretrained models with general architectures. This can provide deeper insights into effective model design and performance. Moreover, Protap introduces two novel specialized applications: enzyme-catalyzed protein cleavage site prediction and targeted protein degradation by PROTACs, which are biological processes not considered by prior benchmarks. A more comprehensive comparison between existing benchmarks is presented in Tab. 1.
3 Protein Modeling in Protap
In this section, we present an overview of how proteins are represented, modeled, and utilized within the Protap benchmark. We begin by outlining the fundamental definitions of protein sequences and structures. We then introduce the three pretraining tasks employed in Protap: masked language modeling, multi-view contrastive learning, and protein family prediction, followed by a summary of the corresponding models. Finally, we describe the downstream applications supported by Protap, which include two specialized tasks, namely Enzyme-Catalyzed Protein Cleavage Site Prediction and Targeted Protein Degradation by Proteolysis-Targeting Chimeras, as well as three general tasks: Protein–Ligand Interactions, Protein Function Annotation Prediction, and Mutation Effect Prediction for Protein Optimization.
3.1 Preliminaries of Proteins
@@
3.2 Pretraining Tasks and Pretraining Models
@@
Recently, the pretraining paradigm has achieved remarkable success across text, images, and graphs [30, 31, 32, 33, 34]. To facilitate protein modeling, various pretraining strategies and model architectures have also been investigated [16, 17, 18, 35]. In this work, our Protap conducts a comprehensive analysis of representative protein pretraining tasks and models to systematically understand their capabilities and limitations in downstream applications. The models and pretraining tasks provided in Protap are summarized in Tab. 2 and briefly introduced below.
@@
Pretraining Task 1: Masked Language Modeling (MLM). This pretext task is based on the sequence information. A protein’s amino acid sequence adheres to an inherent grammar that encodes its structural and functional properties [2]. Therefore, inspired by the success of text modeling, masked language modeling has been adopted to train the protein language model [16] to capture patterns underlying residue sequences. Specifically, as Fig. 1(a) (a) shows, MLM aims to fill in missing amino acids in protein sequences given the masked residue sequence of a protein. ProteinBERT [18], ESM2 [17], and various other protein language models have been pretrained using this MLM task.
@@
Pretraining Task 2: Multi-View Contrastive Learning (MVCL). This pretext task is based on the structural information. Protein structure plays a vital role in determining its biological function. Furthermore, the local structures (motifs) within a protein are biologically related [36, 37]. Inspired by this, multi-view contrastive learning (MVCL) has been extended to preserve the representation similarity between the correlated substructures of proteins [19]. As shown in Fig. 1(a) (b), given a protein $\mathcal{P}$, we generate two views $\mathcal{G}_{seq}^{+}$ and $\mathcal{G}_{space}^{+}$ from its amino acid sequence as positive samples. Views from unrelated proteins serve as negative samples. The MVCL objective is to align positive samples while contrasting them with negative samples in the hidden representation space.
Pretraining Task 3: Protein Family Prediction (PFP). Different from the MLM and MVLC, this task leverages auxiliary protein knowledge to introduce functional and structural supervision into protein representation learning. A protein family is a group of evolutionarily related proteins descended from a common ancestor, typically sharing similar three-dimensional structures, functions, and significant sequence similarities [38, 39]. Consequently, the Protein Family Prediction (PFP) task has been utilized to learn structurally contextualized representations [40]. As illustrated in Fig. 1(a) (c), Protap also deploys the protein family prediction to predict the ground-truth family labels $y$ given the sequence $\mathcal{S}$ and the coordinate $\mathcal{C}$.
Pretraining Models. In Protap, we incorporate and evaluate the following three categories of model architectures pretrained by the aforementioned pretext tasks: (i) Sequence-based model: ProteinBERT [18] and ESM2 [17] treat the protein sequences as biological languages and adopt the transformer [8] to extract the sequential patterns. (ii) Structure-based model: Protap also implementes EGNN [11], SE(3) transformer [41], and GVP [10], which are representative equivariant architectures for structural information encoding. (iii) Sequence-structure hybrid model: SaProt [35] enhances the protein language model with explicit structure vocabulary. Additionally, inspired by Transformer-M [15] and Unizyme [22], Protap implements a D-Transformer, which introduces the residue distance matrix into the sequence attention computation as shown in Tab. 2. Protap pretrains ProteinBERT, EGNN, SE3 transformer, GVP, and D-Transformer on each aforementioned pretraining task with UniProt [42], yielding 15 pretraining models. For the ESM-2 and SaProt, we directly adopt their publicly available pretrained weights for protein modeling. More details of pretraining model architectures, pretraining datasets, and other pretraining details are given in Appendix E.
Table 2: Summary of pretraining models and domain models in our Protap. Models highlighted in (*) are trained from scratch, while those in (*) use only publicly available pretrained weights.
Model Input Modalities Pretrain Data #Parameter Objective Source
Pretrain Models
EGNN [11] AA Seq & 3D Coord Swiss-Prot 540k  [42] 10M MLM, MVCL, PFP ICML, 2021
SE(3) Transformer [41] AA Seq & 3D Coord Swiss-Prot 540k  [42] 4M MLM, MVCL, PFP NeurIPS, 2020
GVP [10] AA Seq & 3D Coord Swiss-Prot 540k  [42] 0.2M MLM, MVCL, PFP ICLR, 2021
ProteinBERT [18] AA Seq Swiss-Prot 540k  [42] 72M MLM, MVCL, PFP Bioinformatics, 2022
D-Transformer [15, 22] AA Seq & 3D Coord Swiss-Prot 540k  [42] 3.5M MLM, MVCL, PFP ArXiv, 2025ICLR, 2023
ESM2 [17] AA Seq UR50 70M [42] 650M MLM Science, 2023
SaProt [35] AA Seq & 3D Coord UR50 40M [42] 650M MLM ICLR, 2024
Domain Specific Models
ClipZyme [13] AA Seq & 3D Coord & SMILES 14.8M PCS ICML, 2024
UniZyme [22] AA Seq & 3D Coord Swiss-Prot 11k  [42] 15.5M PCS ArXiv, 2025
DeepProtacs [23] AA Seq & 3D Coord & SMILES 0.1M PROTACs Nature Communications, 2022
ETProtacs [24] AA Seq & 3D Coord & SMILES 5.4M PROTACs Briefings in Bioinformatic, 2025
KDBNet [43] AA Seq & 3D Coord & SMILES 3.4M PLI Nature Machine Intelligence, 2023
MONN [44] AA Seq & 3D Coord 1.7M PLI Cell Systems, 2024
DeepFRI [45] AA Seq & 3D Coord Pfam 10M  [39] 1.8M PFA Nature Communications, 2021
DPFunc [21] AA Seq & 3D Coord & Protein Domain 110M PFA Nature Communications, 2025
(a) Illustration of pretraining tasks in our Protap. (I) Masked Language Modeling is a self-supervised objective designed to recover masked residues in protein sequences; (II) Multi-View Contrastive Learning leverages protein structural information by aligning representations of biologically correlated substructures. Given two views of the same protein; (III) Protein Family Prediction introduces functional and structural supervision by training models to predict family labels based on protein sequences and 3D structures.
@@
(b) Illustration of downstream applications in our Protap. (I) Biological process of the enzyme-catalyzed protein hydrolysis; (II) Biological process of targeted protein degradation by PROTACs, where PROTACs form a ternary complex with the target protein and E3 ligase, leading to target protein ubiquitination and degradation; (III) In protein-ligand interaction, the ligand binds to the protein pocket, blocking interactions with other molecules; (IV) Protein function annotation prediction reveals the biological activities a protein participates in; (V) Mutation effect prediction optimizes protein properties or functions, facilitating protein engineering.
Figure 1: Overview of pre-training and downstream tasks in Protap.
3.3 Applications and Domain Models

### v3

# arXiv:2506.02052v3[q-bio.BM] 02 Mar 2026

@@ Abstract.
Recently, extensive deep learning architectures and pretraining strategies have been explored to support downstream protein applications. Additionally, domain-specific models incorporating biological knowledge have been developed to enhance performance in specialized tasks. In this work, we introduce Protap, a comprehensive benchmark that systematically compares backbone architectures, pretraining strategies, and domain-specific models across diverse and realistic downstream protein applications. Specifically, Protap covers five applications: two general tasks and three novel specialized tasks, i.e., enzyme-catalyzed protein cleavage site prediction and targeted protein degradation, which are industrially relevant yet missing from existing benchmarks. For each application, Protap compares various domain-specific models and general architectures under multiple pretraining settings. Our empirical studies imply that: (i) Though large-scale pretraining encoders achieve great results, they often underperform supervised encoders trained on small downstream training sets. (ii) Incorporating structural information during downstream fine-tuning can match or even outperform protein language models pretrained on large-scale sequence corpora. (iii) Domain-specific biological priors can enhance performance on specialized downstream tasks. Code is publicly available at https://github.com/Trust-App-AI-Lab/protap.
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Preprint; 2026; booktitle: Preprintjournalyear: 2026isbn: 978-1-4503-XXXX-X/26/08
1. Introduction
@@
Proteins serve as the central executors of biological activities, regulating a wide range of critical biological processes through their complex three-dimensional structures and dynamic properties. A precise understanding of protein function and interactions is critical across many applications (Wong et al., 2024; Anfinsen, 1973). For instance, directed evolution of enzymes can endow proteins with novel functions (Austin et al., 2018). Accurate prediction of protein–ligand interactions (PLIs) can largely accelerate drug discovery (Sadybekov et al., 2022; Zhang et al., 2025; Wu et al., 2024). These application areas underscore the immense potential of deep learning in protein analysis.
@@
Various deep model architectures and pretraining tasks have been developed to leverage protein sequences and structures for protein applications. (i) Geneal Architectures: Sequence models such as LSTM (Hochreiter and Schmidhuber, 1997) and Transformer (Vaswani et al., 2017) have been employed to extract amino acid sequence patterns for tasks like stability landscape prediction (Rao et al., 2019). Geometric Graph Neural Networks (GNNs), including GVP-GNN (Jing et al., 2021) and EGNN (Satorras et al., 2021), have demonstrated remarkable effectiveness in modeling 3D protein structures (Dauparas et al., 2022; Mikhael et al., 2024). Recently, Graphformer (Ying et al., 2021) and Transformer-M (Luo et al., 2023a) further incorporated structural biases into transformer-based architectures. These emerging sequence-structure hybrid models show promising ability in protein representation learning. (ii) Pretrainig Strategies: Recently, large-scale protein pretraining models have also been explored to benefit the representation learning for downstream applications. In particular, masked language modeling has been extended to predict the masked amino acids in protein sequences, resulting in protein language models like ESM and ProteinBERT (Rives et al., 2021; Hayes et al., 2025; Brandes et al., 2022). Morever, GearNet (Zhang et al., 2023) explores the multi-view contrastive learning on protein structures. Additionally, some methods, such as OntoProtein (Zhang et al., 2022), leverage functional annotations as supervision signals for protein encoder pretraining.
@@
Apart from these neural architectures and pretraining tasks for protein modeling, significant progress has also been made in developing domain-specific models for various downstream protein applications. For instance, protein function prediction models such as DPFunc (Wang et al., 2025) leverage sequence, structural, and domain-level information to enhance the accuracy of predicting the Gene Ontology annotations of proteins. Additionally, to improve enzyme modeling, UniZyme (Li et al., 2025) integrates the energy frustration matrix and enzyme active-site knowledge into a transformer-based framework. In the domain of proteolysis-targeting chimera (PROTAC) modeling, dedicated methods such as DeepProtacs (Li et al., 2022) and ETProtacs (Cai et al., 2025) have been developed to accurately capture the ternary interactions essential for targeted protein degradation.
A short description of the image A long description of the image content, detailing all important visual information for a reader using screen-reading software.
Figure 1. Overview of our benchmark Protap.
@@
Given the extensive variety of general model architectures, pretraining strategies, and domain-specific models, there remains a gap in benchmarking general pretrained models alongside domain-specific models. This leads us to a fundamental question: Can large-scale general pretrained models outperform domain-specific models on downstream protein tasks? If so, under what conditions, and which model architectures or pretraining strategies enable such advantages? However, current benchmarks predominantly focus on specific pretrained model categories, such as protein language models in (Xu et al., 2022; Capel et al., 2022) and geometric GNNs in (Jamasb et al., 2024). The comparisons of our Protap with existing benchmarks are given in Appendix Tab. 5. In summary, existing works lack a comprehensive evaluation of general protein model architectures, pretraining strategies, and domain-specific models across realistic biological applications. To address these gaps, we introduce Protap, a standardized benchmark that compares architectures, pre-training strategies, and domain models on five realistic downstream applications. Our contributions are:
@@
  • We identify and integrate realistic applications from existing literature and databases to support comprehensive evaluations. In addition to two general applications, Protap introduces two novel specialized applications: enzyme-catalyzed protein cleavage site prediction and targeted protein degradation by PROTACs, which are biological processes not covered by prior benchmarks. Furthermore, Protein-Ligand Interaction is also categorized as a specialized application due to its complex molecular interactions.
@@
  • Comprehensive and diverse proteins are covered in the benchmark, which include enzymes, receptors, drugs, etc. The tasks are diverse, which include single protein modeling, interaction modeling such as PLI, enzyme-substrate modeling, and complex interaction process modeling, i.e., PROTACs.
@@
  • We compare 18 protein pretraining models and 8 domain models across 5 realistic protein applications. This offers insights into the development of protein foundation models and the design of domain-specific models for downstream applications.
2. Related Works
@@
Existing benchmarks for protein modeling primarily emphasize sequence-based models. TAPE (Rao et al., 2019) introduces standardized sequence-centric tasks, including structure prediction and remote homology detection. PEER (Xu et al., 2022) expands the sequence-based scope to cover protein–protein interactions and functional annotations. ProteinGLUE (Capel et al., 2022) is another sequence-based benchmark designed to evaluate sequence pretraining methods. ProteinGym (Notin et al., 2023) establishes an evaluation framework for mutation effect prediction with protein language models. ProteinBench (YE et al., 2025) evaluates protein foundation models mainly in protein generation tasks. Among existing benchmarks, ProteinWorkshop (Jamasb et al., 2024) incorporates the equivariant graph neural networks to evaluate the structure-based models. However, ProteinWorkshop lacks a unified benchmarking pipeline that systematically compares sequence-based, structure-based, and hybrid models within a standardized framework.
@@
Furthermore, existing benchmarks predominantly focus on general model architectures such as LSTMs (Hochreiter and Schmidhuber, 1997), Transformers (Vaswani et al., 2017), and EGNNs (Satorras et al., 2021). In contrast, Protap explores the advantages of domain-specific model designs and the integration of domain knowledge. Our Protap systematically benchmarks domain-specific models alongside pretrained models with general architectures. This can provide deeper insights into effective model design and performance. Moreover, Protap introduces two novel specialized applications: enzyme-catalyzed protein cleavage site prediction and targeted protein degradation by PROTACs, which are biological processes not considered by prior benchmarks. A more comprehensive comparison between existing benchmarks is presented in Tab. 5.
Table 1. Overview of models and datasets used in Protap.
(a) Pretraining models and domain models in our Protap. Models highlighted in (*) are trained from scratch, while those in (*) use only publicly available pretrained weights.
Model Input Modalities Pretrain Data #Parameter Training Objective Source
Pretrain Models
EGNN AA Seq & 3D Coord Swiss-Prot 540k 10M MLM, MVCL, PFP ICML, 2021
SE(3) Transformer AA Seq & 3D Coord Swiss-Prot 540k 4M MLM, MVCL, PFP NeurIPS, 2020
GVP AA Seq & 3D Coord Swiss-Prot 540k 2M MLM, MVCL, PFP ICLR, 2021
ProteinBERT AA Seq Swiss-Prot 540k 72M MLM, MVCL, PFP Bioinformatics, 2022
D-Transformer AA Seq & 3D Coord Swiss-Prot 540k 3.5M MLM, MVCL, PFP NeurIPS, 2025ICLR, 2023
ESM-2 AA Seq UR50 70M 650M MLM Science, 2023
ESM Cambrian AA Seq UR70, MGnify, JGI 3B 600M MLM Science, 2025
SaProt AA Seq & 3D Coord UR50 40M 650M MLM ICLR, 2024
Domain Specific Models
ClipZyme AA Seq & 3D Coord & SMILES 14.8M PCS ICML, 2024
UniZyme AA Seq & 3D Coord Swiss-Prot 11k 15.5M PCS NeurIPS, 2025
DeepProtacs AA Seq & 3D Coord & SMILES 0.1M PROTACs Nature Communications, 2022
ETProtacs AA Seq & 3D Coord & SMILES 5.4M PROTACs Briefings in Bioinformatic, 2025
KDBNet AA Seq & 3D Coord & SMILES 3.4M PLI Nature Machine Intelligence, 2023
MONN AA Seq & 3D Coord 1.7M PLI Cell Systems, 2024
DeepFRI AA Seq & 3D Coord Pfam 10M 1.8M PFA Nature Communications, 2021
DPFunc AA Seq & 3D Coord & Protein Domain 110M PFA Nature Communications, 2025
(b) Summary of datasets and metrics for various prediction tasks.
Application Category Data Source #Train #Test Metric
Pretraining (Consortium, 2019) 542,378
Protein Cleavage Site Prediction Specialized (Rawlings et al., 2014) 1609 316 AUC, AUPR
Targeted Protein Degradation Specialized (Ge et al., 2025) 843 209 Acc, AUC
Protein–Ligand Interactions Specialized (Luo et al., 2023b) 11,520 2,880 MSE, Pearson
Protein Function Annotation Prediction General (Gligorijević et al., 2021; Jamasb et al., 2024) 23,760 2,023 Fmax, AUPR
Mutation Effect Prediction General (Notin et al., 2023) 2.4M AUC, Pearson
Refer to caption A short description of the image A long description of the image content, detailing all important visual information for a reader using screen-reading software.
Figure 2. Illustration of pretraining tasks in Protap.
3. Protein Pretraining Models in Protap
In this section, we first outline the fundamental definitions of protein sequences and structures. We then introduce the three pretraining tasks employed in Protap: masked language modeling, multi-view contrastive learning, and protein family prediction. Finally, we summarize the protein pretraining models deployed in the Protap benchmark.
3.1. Preliminaries of Proteins
@@
3.2. Pretraining Tasks and Pretraining Models
@@
Recently, the pretraining paradigm has achieved remarkable success across text, images, and graphs (Devlin et al., 2019; Achiam et al., 2023; Radford et al., 2021; Hu et al., 2020). To facilitate protein modeling, various pretraining strategies and model architectures have also been investigated (Rives et al., 2021; Lin et al., 2023; Brandes et al., 2022; Su et al., 2024). In this work, our Protap conducts a comprehensive analysis of representative protein pretraining tasks and models to systematically understand their capabilities and limitations in downstream applications. The models and pretraining tasks provided in Protap are summarized in Tab. 1 and briefly introduced below.
Pretraining Task 1: Masked Language Modeling (MLM).
@@
This pretext task is based on the sequence information. A protein’s amino acid sequence adheres to an inherent grammar that encodes its structural and functional properties (Anfinsen, 1973). Therefore, inspired by the success of text modeling, masked language modeling has been adopted to train the protein language model (Rives et al., 2021) to capture patterns underlying residue sequences. Specifically, as Fig. 2 (a) shows, MLM aims to fill in missing amino acids in protein sequences given the masked residue sequence of a protein. ProteinBERT (Brandes et al., 2022), ESM2 (Lin et al., 2023), ESM Carbrain (Hayes et al., 2025), and various other protein language models have been pretrained using this MLM task.
Pretraining Task 2: Multi-View Contrastive Learning (MVCL).
@@
This pretext task is based on the structural information. Protein structure plays a vital role in determining its biological function. Furthermore, the local structures (motifs) within a protein are biologically related (Ponting and Russell, 2002; Mackenzie and Grigoryan, 2017). Inspired by this, multi-view contrastive learning (MVCL) has been extended to preserve the representation similarity between the correlated substructures of proteins (Zhang et al., 2023). As shown in Fig. 2 (b), given a protein $\mathcal{P}$, we generate two views $\mathcal{G}_{seq}^{+}$ and $\mathcal{G}_{space}^{+}$ from its amino acid sequence as positive samples. Views from unrelated proteins serve as negative samples. The MVCL objective is to align positive samples while contrasting them with negative samples in the hidden representation space.
A short description of the image A long description of the image content, detailing all important visual information for a reader using screen-reading software.
@@
Figure 3. Illustration of downstream applications in Protap. (a) Biological process of the enzyme-catalyzed protein hydrolysis; (b) Biological process of targeted protein degradation by PROTACs, where PROTACs form a ternary complex with the target protein and E3 ligase, leading to target protein ubiquitination and degradation; (c) In protein-ligand interaction, the ligand binds to the protein pocket, blocking interactions with other molecules; (d) Protein function annotation prediction reveals the biological activities a protein participates in; (e) Mutation effect prediction optimizes protein properties or functions for protein engineering.
Pretraining Task 3: Protein Family Prediction (PFP).
Different from the MLM and MVLC, this task leverages auxiliary protein knowledge to introduce functional and structural supervision into protein representation learning. A protein family is a group of evolutionarily related proteins descended from a common ancestor, typically sharing similar three-dimensional structures, functions, and significant sequence similarities (Sillitoe et al., 2021; Bateman et al., 2004). Consequently, the Protein Family Prediction (PFP) task has been widely utilized to learn structurally contextualized representations (Min et al., 2021). As illustrated in Fig. 2 (c), Protap also deploys the protein family prediction to predict the ground-truth family labels $y$ given the sequence $\mathcal{S}$ and the associated coordinate $\mathcal{C}$.
3.3. Pretraining Models
In Protap, we incorporate the following three categories of model architectures pretrained by the aforementioned pretext tasks:
  • Sequence-based model: ProteinBERT (Brandes et al., 2022), ESM-2 (Lin et al., 2023) and ESM Carbrain (Hayes et al., 2025) treat the protein sequences as biological languages and adopt the transformer (Vaswani et al., 2017) to extract the sequential patterns.
  • Structure-based model: Protap also implementes EGNN (Satorras et al., 2021), SE(3) transformer (Fuchs et al., 2020), and GVP (Jing et al., 2021), which are representative equivariant architectures for structural information encoding.
  • Sequence-structure hybrid model: SaProt (Su et al., 2024) enhances the protein language model with explicit structure vocabulary. Additionally, inspired by Transformer-M (Luo et al., 2023a) and Unizyme (Li et al., 2025), Protap implements a D-Transformer, which introduces the residue distance matrix into the sequence attention computation as shown in Tab. 1.
Protap pretrains ProteinBERT, EGNN, SE3 transformer, GVP, and D-Transformer on each pretraining task with UniProt (Consortium, 2019), yielding 15 pretraining models. For the ESM-2, ESM-C and SaProt, we directly adopt their publicly available pretrained weights for protein modeling. More details of pretraining model architectures, pretraining datasets, and other pretraining details are given in Appendix G.
4. Applications and Domain Models